kekeman:

A circle is centered at (7, 8) and has a radius of 11. Which of the following is the equation for this circle? (x − 7)2 + (y − 8)2 = 121 (x − 7)2 + (y − 8)2 = 11 (x + 7)2 + (y + 8)2 = 121 (x + 7)2 + (y + 8)2 = 11

wellsh24:

Ok so does anyone have an essay on why internet access should be free i really need it by the end of the day today

Florisalreadytaken:

so we have this formula: \[ (x− a)2 + (y−b )2 = r^2 ​\] \[(a,b) \Rightarrow (7,8)\] \[ r=11 \] plugging that info in the formula, it would be: \[ (x−7)2 + (y−8 )2 = 11^2 \] \[ \updownarrow \] \[ (x−7)2 + (y−8 )2 = 121 \]

Florisalreadytaken:

that said, which option is the right one?

kekeman:

It is going to be A

Florisalreadytaken:

there you go!